### v1

# Cluster-Based Normalization Layer for Neural Networks

Abstract
Deep learning faces significant challenges during the training of neural networks, including internal covariate shift, label shift, vanishing/exploding gradients, overfitting, and computational complexity. While conventional normalization methods, such as Batch Normalization, aim to tackle some of these issues, they often depend on assumptions that constrain their adaptability. Mixture Normalization faces computational hurdles in its pursuit of handling multiple Gaussian distributions. This paper introduces Cluster-Based Normalization (CB-Norm) in two variants—Supervised Cluster-Based Normalization (SCB-Norm) and Unsupervised Cluster-Based Normalization (UCB-Norm)—introducing a groundbreaking one-step normalization approach. CB-Norm leverages a Gaussian mixture model to specifically address challenges related to gradient stability and learning acceleration. For SCB-Norm, a supervised variant, the novel mechanism involves introducing predefined data partitioning, termed clusters, to normalize activations based on the assigned cluster. This cluster-driven approach creates a space that conforms to a Gaussian mixture model. On the other hand, UCB-Norm, an unsupervised counterpart, dynamically clusters neuron activations during training, adapting to task-specific challenges without relying on predefined data partitions (clusters). This dual approach ensures flexibility in addressing diverse learning scenarios. CB-Norm innovatively uses a one-step normalization approach, where parameters of each mixture component (cluster in activation space) serve as weights for deep neural networks. This adaptive clustering process tackles both clustering and resolution of deep neural network tasks concurrently during training, signifying a notable advancement in the field.
I Introduction
Normalization in the context of deep learning is a fundamental data transformation process aimed at conferring specific statistical properties upon the data. Within the domain of data normalization, various techniques are employed, such as centering, scaling, standardizing, decorrelating, and whitening [1]. In deep neural networks (DNNs), input normalization plays a crucial role during training by eliminating the disparities in feature magnitudes. This normalization process facilitates the rapid convergence of linear model optimization [2]. However, in complex layered neural networks, where the input is directly linked only to the initial weight matrix, it becomes less transparent how the input influences the optimization landscape in relation to other weight matrices. Typically, these initial weights are not normalized, and this omission can significantly impact the gradient optimization procedure. To address this issue, various weight-initialization techniques have been introduced, aiming to equalize variances in layer input/output gradients across different layers [3]. Nevertheless, as weight matrices are updated throughout training, the equal variance property may become disrupted. From this perspective, the normalization of activations during training emerges as a pivotal strategy to maintain the advantages associated with normalizing inputs. By ensuring that activations maintain a consistent statistical distribution across layers, deep neural networks can benefit from stable and efficient training, ultimately leading to improved model performance.
Batch Normalization (BN), as introduced by Ioffe and Szegedy in their seminal work [4], stands out as the prevailing and widely adopted method for the normalization of activations. BN achieves this by standardizing activations using batch statistics, which facilitates the utilization of higher learning rates. Nevertheless, BN exhibits limitations tied to batch size dependence and the assumption of uniform data distribution. To mitigate these limitations, specialized variants of BN have been developed [5], focusing on addressing issues related to batch size. On a different front, Mixture Normalization (MN) [6] offers an alternative approach, aiming to overcome the constraints imposed by data distribution assumptions. In stark contrast to BN, MN accommodates data samples drawn from diverse distributions. It employs the Expectation-Maximization (EM) algorithm [7] for data clustering based on components, allowing for the normalization of each sample using component-specific statistics. MN’s distinctive approach enhances convergence, particularly in convolutional neural networks, when compared to BN. However, the incorporation of the EM algorithm in mixture models may entail a notable increase in computational time, potentially hampering overall efficiency. This stems from the necessity of estimating parameters across the entire dataset prior to neural network training, ensuring the provision of normalization parameters.
To address this issues, we propose a new normalization method titled Cluster-Based Normalization (CB-Norm). CB-Norm serves as a normalization layer within deep neural network architectures. Operating on the hypothesis that activations can be represented as a Gaussian mixture model, CB-Norm normalizes these activations during deep neural network training to estimate parameters for each mixture component. These parameters are treated as learnable parameters and are updated as neural network weights during backpropagation, aligning with the target task. To operationalize this strategy, we suggest two approaches: a supervised version (SCB-Norm) and an unsupervised version (UCB-Norm).
SCB-Norm operates in two steps: first, creating clusters to group samples with similar characteristics, and then normalizing activations during training based on these clusters (activations from samples within the same cluster are normalized with the same parameters). Building clusters involves expert knowledge for effective data grouping, or it can be derived from partitioning using another clustering algorithm. For example, in a dataset with classes and superclasses (groupings of classes), each superclass can be treated as a distinct cluster. In domain adaptation, each specific domain within the dataset can be considered an independent cluster.
UCB-Norm is a one-step method where the normalization of activations during neural network training is not conditioned on prior knowledge (clusters). The mixture components in the activation space are constructed by considering the contribution of all parameters of each mixture component for the normalization of activations.
It’s worth noting that the choice of a Gaussian distribution assumption for clusters is made to ensure a coherent basis for comparison. However, it is important to emphasize that alternative hypotheses for data distribution could also be considered in different scenarios and experiments. As an integral layer within the deep neural network, CB-Norm plays a pivotal role in standardizing activations originating from the same cluster using the parameters acquired through backpropagation. This process facilitates the estimation of parameters for each cluster, ultimately enhancing the data representation’s discriminative capacity with respect to the target task.
This study presents four noteworthy contributions:
  • Cluster-Based Normalization (CB-Norm) Advancements: CB-Norm introduces an approach dynamically learning mixture component parameters during neural network training. Notably, CB-Norm excels in gradient stability, learning acceleration, and adaptability across diverse learning scenarios.
  • Two CB-Norm Applications for Enhanced Normalization: We present two CB-Norm methods - Supervised Cluster-Based Normalization (SCB-Norm), utilizing predefined data partitioning based (clusters), and Unsupervised Cluster-Based Normalization (UCB-Norm), which dynamically clusters neuron activations during training for increased flexibility.
  • Versatile Application in Deep Learning Architectures: CB-Norm exhibits versatility across various deep neural network architectures, including Transformers and Convolutional Neural Networks. It serves as a layer at different depths, expediting training processes and improving generalization performance consistently.
  • Integration Recommendations for Improved Performance: We recommend integrating CB-Norm into domains such as domain adaptation and Generative Adversarial Networks (GANs). CB-Norm addresses challenges in domain adaptation effectively, enhancing model performance across diverse data distributions. Moreover, CB-Norm contributes to resolving issues like ”mode collapse” in GANs, improving their learning from a diverse set of modes.
II Related Work
II-A Batch Normalization
To ensure clear understanding and consistency in our model, and to facilitate comparison with existing work, we adopt the same notations as those used in [6]. In a convolutional neural network layer, let $x\in\mathbb{R}^{N\times C\times H\times W}$ represent the neuron’s activation, where $N$, $C$, $H$, and $W$ denote batch, channel, height, and width axes. Batch normalization (BN) [4] standardizes $x$ with $m$ samples mini-batch $B=\{x_{1:m}:m\in[1,N]\times[1,H]\times[1,W]\}$, flattening $x$ except the channel axis, as follows:
$\hat{x}_{i}=\frac{x_{i}-\mu_{B}}{\sqrt{\sigma^{2}_{B}+\epsilon}},$ (1)
where $\mu_{B}=\frac{1}{m}\sum_{i=1}^{m}x_{i}$ and $\sigma^{2}_{B}=\frac{1}{m}\sum_{i=1}^{m}(x_{i}-\mu_{B})^{2}$ represent the mean and variance, while $\epsilon>0$ serves as a small value to mitigate numerical instability. To mitigate the constraints imposed by standardization, BN introduces learnable parameters $\gamma$ and $\beta$ to remove the linear regime of nonlinearity in certain activations:
$\tilde{x}_{i}=\gamma\hat{x}_{i}+\beta$
When the samples in the mini-batch originate from the same distribution, the transformation described in Equation 1 yields a distribution with a mean of zero and a variance of one. This constraint of zero mean and unit variance serves to stabilize the activation distribution, ultimately aiding in the training process.
While BN demonstrates effective performance, it exhibits two notable limitations: its reliance on batch size, leading to suboptimal performance with small batches, and the simplistic assumption that sample within the mini-batch are from the same distribution, a scenario rarely encountered in real-world data. In response to these challenges, several alternative approaches have been introduced.
II-B Extensions of Batch Normalization
Several extensions to batch normalization (BN) have been introduced, including Layer Normalization (LN) [8], Instance Normalization (IN) [9], Group Normalization (GN) [10], and Mixture Normalization (MN) [6]. These variants all employ the common transformation $x\rightarrow\hat{x}$, which is applied to the flattened $x$ along the spatial dimension $L=H\times W$, as follows:
$v_{i}=x_{i}-\mathbb{E}_{B_{i}}(x),\ \hat{x}_{i}=\frac{v_{i}}{\sqrt{\mathbb{E}_{B_{i}}(v^{2})+\epsilon}},$ (2)
where $B_{i}=\{j:j_{N}\in[1,N],j_{C}\in[i_{C}],j_{L}\in[1,L]\}$, and $i=(i_{N},i_{C},i_{L})$ a vector indexing the activations $x\in\mathbb{R}^{N\times C\times L}$.
Layer Normalization (LN): Alleviates inter-dependency among batch-normalized activations by computing mean and variance using input from individual layer neurons. LN proves beneficial for recurrent networks but can encounter difficulties with convolutional layers due to spatial variations in visual data. It can be formulated as Equation 2 when $B_{i}=\{j:j_{N}\in[i_{N}],j_{C}\in[1,C],j_{L}\in[1,L]\}$.
Instance Normalization: A normalization approach that normalizes each sample individually, primarily aimed at eliminating style information, particularly in images. IN enhances the performance of specific deep neural networks (DNNs) by computing mean values and standard deviations in the spatial domain. It finds extensive use in applications such as image style transfer [11] [12]. It can be formulated as Equation 2 when $B_{i}=\{j:j_{N}\in[i_{N}],j_{C}\in[i_{C}],j_{L}\in[1,L]\}$.
Group Normalization (GN): A normalization method that partitions neurons into groups and autonomously standardizes layer inputs for each sample within these groups. GN performs exceptionally well in visual tasks characterized by small batch sizes, such as object detection and segmentation. GN partitions the channels into several groups ($G$), calculating statistics along the $L$ axis, but restricting this computation to subgroups of the channels. When $G=C$, GN is equivalent to IN, and it is identical to LN when $G=1$.
Mixture Normalization (MN): In the realm of deep neural networks (DNNs), the distribution of activations inherently encompasses multiple modes of variation due to the presence of non-linearities. The batch normalization (BN) [4] assumption that a Gaussian distribution suffices to model the generative process of mini-batch samples becomes less applicable. In response, Mixture Normalization (MN) [6] approaches BN from the perspective of Fisher kernels, originating from generative probability models. Instead of computing mean and standard deviation across entire instances within a mini-batch, MN employs a Gaussian Mixture Model (GMM) to assign each instance in the mini-batch to a component and subsequently normalizes with regard to multiple means and standard deviations associated with different modes of variation in the underlying data distribution.
MN normalizes each sample within the mini-batch by utilizing the mean and standard deviation corresponding to the mixture component to which the sample is assigned. The probability density function $p_{\theta}$ that characterizes the data can be represented as a parameterized Gaussian Mixture Model (GMM). Let $x\in\mathbb{R}^{D}$, and $\theta=\{\lambda_{k},\mu_{k},\Sigma_{k}:k=1,...,K\}$,
$p(x)=\sum_{k=1}^{K}\lambda_{k}p(x|k),\ s.t.\ \forall_{k}\ :\ \lambda_{k}\ \geq 0,\ \sum_{k=1}^{K}\lambda_{k}=1,$ (3)
where
$p(x|k)=\frac{1}{(2\pi)^{D/2}\lvert\Sigma_{k}\rvert^{1/2}}\exp\left(-\frac{(x-\mu_{k})^{T}\Sigma_{k}^{-1}(x-\mu_{k})}{2}\right),$ (4)
is the $k^{th}$ Gaussian in the mixture model $p(x)$, $\mu_{k}$ the mean vector and $\Sigma_{k}$ is the covariance matrix.
The probability that $x$ has been generated by the $k^{th}$ Gaussian component in the mixture model can be defined as:
$\tau_{k}(x)=p(k|x)=\frac{\lambda_{k}p(x|k)}{\sum_{j=1}^{K}\lambda_{j}p(x|j)},$ (5)
Based on these assumptions and the general transform in Equation (2), the Mixture Normalizing Transform for a given $x_{i}$ is defined as
$\hat{x}_{i}=\sum_{k=1}^{K}\frac{\tau_{k}(x_{i})}{\sqrt{\lambda_{k}}}\hat{x}_{i}^{k},$ (6)
given
$v_{i}^{k}=x_{i}-\mathbb{E}_{B_{i}}[\hat{\tau}_{k}(x).x],\ \hat{x}_{i}^{k}=\frac{v_{i}^{k}}{\sqrt{\mathbb{E}_{B_{i}}[\hat{\tau}_{k}(x).(v^{k})^{2}]+\epsilon}},$ (7)
where
$\hat{\tau}_{k}(x_{i})=\frac{\tau_{k}(x_{i})}{\sum_{j\in B_{i}}\tau_{k}(x_{j})},$
is the normalized contribution of $x_{i}$ with respect to the mini-batch $B_{i}=\{j:j_{N}\in[i_{N}],j_{C}\in[i_{C}],j_{L}\in[1,L]\}$ in the estimation of the statistical measures of the $k^{th}$ Gaussian component.
With this approach, Mixture Normalization can be implemented in two stages:
  • Estimation of the mixture model’s parameters $\theta=\{\lambda_{k},\mu_{k},\Sigma_{k}:k=1,\ldots,K\}$ using the Expectation-Maximization (EM) algorithm [7].
  • Normalization of each sample based on the estimated parameters (see Equation (7)) and aggregation using posterior probabilities (see Equation (6)).
In the context of convolutional neural networks, this method enables Mixture Normalization to outperform batch normalization in terms of convergence and accuracy in supervised learning tasks.
III Proposed Method: Cluster-Based Normalization
Our Cluster-Based Normalization (CB-Norm) method introduces a one-step normalization method, simultaneously clustering neuron activations and addressing deep neural network tasks. By leveraging a latent space assumption with neuron activations following a Gaussian mixture model, AN learns parameters for each mixture component during neural network training. These learned parameters serve as weights, contributing to adaptive components tailored to the target task, thereby enhancing gradient stability, accelerating convergence, and improving overall neural network performance. The implementations of CB-Norm, namely Supervised Cluster-Based Normalization (SCB-Norm) and Unsupervised Cluster-Based Normalization (UCB-Norm), provide specific strategies for incorporating cluster information into the normalization process.
III-A Supervised Cluster-Based Normalization (SCB-Norm)
In SCB-Norm, we leverage prior knowledge called cluster. A cluster is simply a grouping of data that shares similar characteristics. Creating clusters requires either specialized expertise to effectively group data or can be accomplished through partitioning using an alternative clustering algorithm. For instance, in a dataset containing classes and superclasses (collections of classes), each superclass can be regarded as a separate cluster. In the context of domain adaptation, each distinct domain within the dataset can be treated as an independent cluster. SCB-Norm uses these clusters to condition the construction of the Gaussian Mixture Model (GMM). This implies that activations from samples within the same cluster undergo normalization using identical parameters, which are learned as part of the deep neural network and are updated during backpropagation based on the target task. On the latent space of activations, clusters are expected to follow a GMM. SCB-Norm essentially tailors the normalization process based on predefined clusters, enhancing adaptability to specific features in the data.
Let $x$ denote an activation tensor in the $\mathbb{R}^{N\times C\times H\times W}$ space. Here, $B_{i}=\{j:j_{N}\in[i_{N}],j_{C}\in[i_{C}],j_{L}\in[1,L]\}$ signifies a set of activations obtained by flattening $x$ along the $L=H\times W$ axis. SCB-Norm functions by standardizing each activation $x_{i}$ in $B_{i}$, derived from the transformation of a sample associated with a specific mixture component $k$. The normalization is carried out using mixture component-specific parameters $\theta_{k_{i}}=\{\mu_{k_{i}},\sigma_{k_{i}},\lambda_{k_{i}}\}$, where $\lambda_{k_{i}}$ represents a constant, assuming that mixture components are equiprobable. This assumption implies that each mixture component contributes equally to the overall distribution, simplifying the weighting scheme within the Gaussian Mixture Model (GMM):
$SCB-Norm_{\theta_{k_{i}}}(x_{i}):\hat{x}_{i}\leftarrow\frac{x_{i}-\mu_{k_{i}}}{\sqrt{\sigma^{2}_{k_{i}}+\epsilon}}$ (8)
In the backpropagation process, the deep neural network dynamically acquires parameters $\theta=\{\mu_{k},\sigma_{k}\}_{k=1}^{K}$ as weights of the neural network (see Figure 1), with $K$ representing the number of specified mixture components. Notably, $\lambda_{k}$, being a constant, remains unchanged. The continuous learning through backpropagation ensures adaptive updates that align the deep neural network representations with the nuances of the target task. This dynamic adaptation contributes to an iterative refinement of the model, allowing it to capture intricate patterns and variations in the data.
Figure 1: Supervised Cluster-Based Normalization (SCB-Norm) applied on activation $x_{i}$. The mixture component identifier $k_{i}$ determines the cluster to which $x_{i}$ belongs, facilitating the selection of the respective trainable parameters $\theta_{k_{i}}=\{\mu_{k_{i}},\sigma_{k_{i}}\}$ for normalization. The normalized activation, as computed using Equation 8, serves as input for the Deep Neural Network (DNN). Subsequent to loss computation, the parameter $\theta_{k_{i}}$ undergo updates based on the target task.
Each normalized activation $\hat{x}_{i}$ can be viewed as an input to a sub-network composed of the linear transform (Equation 8), followed by the other processing done by the original network. The processing within this sub-network plays a crucial role in amplifying feature extraction, thereby enriching the overall representational capacity of the model. During the training process, as outlined in Algorithm 1, it is essential to backpropagate the gradient of the loss function $\ell$ through this transformation.
1
Input : Deep neural network $Net$ with trainable parameters $\Theta$; subset of activations and its clusters $\{x_{i},k_{i}\}_{i=1}^{m}$, with $k_{i}\in\{1,...,K\}$, where $K$ is the number of clusters
Output : Supervised Cluster-based Normalized deep neural network for inference, $Net^{inf}_{SCB-Norm}$
2 $Net^{tr}_{SCB-Norm}=Net$ // Training SCB-Norm deep neural network
3 for $i\leftarrow 1$ to $m$ do
4
  • Add transformation $\hat{x}_{i}$ = $SCB-Norm_{\theta_{k_{i}}}$($x_{i}$) to $Net^{tr}_{SCB-Norm}$ (Equation 8)
  • Modify each layer in $Net^{tr}_{SCB-Norm}$ with input $x_{i}$ to take $\hat{x}_{i}$ instead
5 end for
6Train $Net^{tr}_{SCB-Norm}$ to optimize the parameters: $\Theta=\Theta\cup\{\mu_{k},\sigma_{k}\}_{k=1}^{K}$
7 $Net^{inf}_{SCB-Norm}=Net^{tr}_{SCB-Norm}$ // Inference SCB-Norm network with frozen parameters
8 for $i\leftarrow 1$ to $m$ do
9
  • Retrieve the parameters associated with the cluster of $x_{i}$: $\theta_{k_{i}}$
  • transform $\hat{x}_{i}$ = $SCB-Norm_{\theta_{k_{i}}}$($x_{i}$) using Equation 8
10 end for
Algorithm 1 Training an Supervised Adaptive-Normalized Network
Moreover, it is imperative to compute gradients with respect to the SCB-Norm transformation’s parameters. This gradient computation involves employing the chain rule, as demonstrated in the forthcoming expression (prior to any simplifications):
$\frac{\partial\ell}{\partial\mu_{k_{i}}}=\frac{\partial\ell}{\partial\hat{x}_{i}}.\frac{\partial\hat{x}_{i}}{\partial\mu_{k_{i}}}=-\frac{\partial\ell}{\partial\hat{x}_{i}}.(\sigma_{k_{i}}^{2}+\epsilon)^{-1/2}$
$\frac{\partial\ell}{\partial\sigma_{k_{i}}^{2}}=\frac{\partial\ell}{\partial\hat{x}_{i}}.\frac{\partial\hat{x}_{i}}{\partial\sigma_{k_{i}}^{2}}=\frac{\mu_{k_{i}}+x_{i}}{2(\sigma_{k_{i}}^{2}+\epsilon)^{3/2}}$
SCB-Norm serves as a differentiable operation within deep neural networks, enabling activations normalization. The core of SCB-Norm involves the construction of a GMM in the latent space. This GMM is defined by parameters $\theta=\{\mu_{k},\sigma_{k}\}_{k=1}^{K}$, which are learned during training. It continually adapts model representations to the target task, enhancing overall performance. The normalization process effectively mitigates variations in activation distributions, allowing the model to emphasize relevant patterns and features. By introducing the GMM in the latent space, SCB-Norm captures intricate relationships in the data. The differentiability of SCB-Norm ensures efficient gradient flow during training, aiding parameter updates while preserving differentiation through the normalization process. This adaptability in the latent space facilitates improved model generalization and task-specific adaptation.
In domain adaptation, where each domain acts as a distinct cluster, the normalization applied to each cluster yields localized, task-specific representations. This approach positively impacts data representation for both source and target domains, aligning them with the target task. SCB-Norm, as a differentiable operation in deep neural networks, underpins this adaptation by continuously fine-tuning model representations for the specific task, ultimately boosting overall performance.
III-B Unsupervised Cluster-Based Normalization (UCB-Norm)
In UCB-Norm, we take a different approach by not relying on prior knowledge (clusters). Instead, the deep neural network autonomously discovers a latent space over activations that adheres to a GMM. UCB-Norm dynamically clusters neuron activations during training without predefined partitions. This allows the model to adapt to task-specific challenges without relying on prior cluster information. The absence of predefined partitions provides UCB-Norm with increased flexibility, as it allows the deep neural network to explore and adapt to patterns in the data on its own.
1
Input : Deep neural network $Net$ with trainable parameters $\Theta$; subset of activations $\{x_{i}\}_{i=1}^{m}$; number of clusters $K$; momentum $m$; $update\in\{weight,moving\ average\}$
Output : Unsupervised Cluster-Based Normalized deep neural network for inference, $Net^{inf}_{UCB-Norm}$
2 $Net^{tr}_{UCB-Norm}=Net$ // Training UCB-Norm deep neural network
3 Initialize the parameters for each cluster as follows:
($\lambda_{k}$, $\mu_{k}$, $\sigma^{2}_{k}$) for $k\in\{1,...,K\}$, subject to the conditions that $\sum_{k=1}^{K}\lambda_{k}=1$ and $\sigma^{2}_{k}>0$ for all $k\in\{1,...,K\}$.
4for $i\leftarrow 1$ to $m$ do
5
  • Add transformation $\hat{x}_{i}$ = $UCB-Norm_{\theta}$($x_{i}$) to $Net^{tr}_{UCB-Norm}$ (Equation 6):
    $\hat{x}_{i}\leftarrow\sum_{k=1}^{K}\frac{\tau_{k}(x_{i})}{\sqrt{\lambda_{k}}}\hat{x}_{i}^{k}$
  • Modify each layer in $Net^{tr}_{UCB-Norm}$ with input $x_{i}$ to take $\hat{x}_{i}$ instead
6 end for
7if $update=weight$ then
8       // Update parameters as weights of the neural network
9       Train $Net^{tr}_{UCB-Norm}$ to optimize the parameters: $\Theta=\Theta\cup\{\lambda_{k},\mu_{k},\sigma_{k}\}_{k=1}^{K}$
10
11 end if
12if $update=moving\ average$ then
13       // Update parameters using moving average according the batch
14       for $r\leftarrow 1$ to $K$ do
15             $\lambda_{k}\leftarrow m\times\lambda_{k}+(1-m)\times mean(\tau_{k})$
16             $\mu_{k}\leftarrow m\times\mu_{k}+(1-m)\times mean(\hat{x}^{k})$
17             $\sigma^{2}_{k}\leftarrow m\times\sigma^{2}_{k}+(1-m)\times variance(\hat{x}^{k})$
18       end for
19
20 end if
[…]
22 for $i\leftarrow 1$ to $m$ do
23
  • Transform $\hat{x}_{i}=UCB-Norm_{\theta}(x_{i})$ using Equation 6
    $\hat{x}_{i}\leftarrow\sum_{k=1}^{K}\frac{\tau_{k}(x_{i})}{\sqrt{\lambda_{k}}}\hat{x}_{i}^{k}$
  • Modify each layer in $Net^{inf}_{UCB-Norm}$ with input $x_{i}$ to take $\hat{x}_{i}$ instead
24 end for
Algorithm 2 Training an Unsupervised Cluster-Based Normalization Network
To align with the representation SCB-Norm, UCB-Norm introduces a technical approach that revolves around estimating the parameters of a Gaussian mixture model (GMM) during the normalization process. Let $K$ denote the number of components in the mixture, and $\theta=\{\lambda_{k},\mu_{k},\sigma_{k}\}_{k=1}^{K}$. As part of the normalization procedure, given the absence of prior knowledge regarding the context of each sample, UCB-Norm employs a strategy involving the consideration of all mixture components for the normalization of $x_{i}$, utilizing Equation 6. This method is designed to aggregate the normalization of $x_{i}$, ensuring that each mixture component contributes proportionally to the normalization process (see Figure 2). By doing so, the mixture component to which $x_{i}$ belongs exerts a more substantial influence on the normalization of $x_{i}$. Through this iterative process, activations within a given cluster are projected onto a shared space, following a Gaussian distribution tailored to the specifics of the target task. The beauty of this approach lies in its adaptability, allowing the model to dynamically adjust its representations based on the characteristics of the data and the learning task at hand. Consequently, UCB-Norm’s technical methodology ensures that the normalization of activations is contextually informed, leading to a more refined and task-adaptive Gaussian mixture model.
Figure 2: Unsupervised Cluster-Based Normalization (UCB-Norm) applied on activation $x_{i}$. The parameter $K$ corresponds to the number of mixture components to be estimated during the training process. In the forward pass, the mixture component parameters $\theta=\{\lambda_{k},\mu_{k},\sigma_{k}\}_{k=1}^{K}$ are utilized to normalize the activation $x_{i}$ through Equation 6. The resulting normalized activation, serves as input for the Deep Neural Network (DNN). After loss computation, in the backward pass, the parameters $\theta$ undergo updates based on the target task.
The update of these mixture component parameters involves two approaches (see Algorithm 2): the first corresponds to an update aligned with the target task ensuring alignment with specific objectives. This targeted update strategy enhances the model’s capability to specialize and optimize its performance for the intended task, resulting in improved task-specific learning and generalization. The second approach consists of estimating parameters using the moving averages method within each batch. This method offers several benefits, including enhanced stability and robustness in the learning process. By smoothing out fluctuations in parameter values across batches, the moving averages method contributes to a more consistent and reliable training trajectory, mitigating the impact of noisy or outlier batches. Additionally, this approach provides a form of regularization, preventing overfitting and promoting better generalization to unseen data. In summary, the dual strategy of updating parameters through both task-aligned updates and the moving averages method combines the benefits of task-specific adaptation and robust, stable learning, contributing to the overall effectiveness and versatility of the training process.
IV Experiments
In the upcoming experiments, we plan to evaluate how well the Cluster-Based Normalization (CB-Norm) performs and to conduct a comparison with Batch Normalization (BN) and Mixture Normalization (MN).
We suggest applying both implementations of CB-Norm: Supervised version (SCB-Norm) and Unsupervised version (UCB-Norm)—to tasks involving classification (Section IV-BIV-C and IV-F), domain adaptation (Section IV-E), and generative neural networks (Section IV-D).
In addition to SCB-Norm, we will incorporate a variant known as SCB-Norm-base. In SCB-Norm-base, two separate neural networks estimate the $K$ clusters (mixture components) parameters $\theta=\{\mu_{k},\sigma_{k}\}_{k=1}^{K}$. For a given input $x_{i}$ in the SCB-Norm-base layer, the cluster identifier $k_{i}$ of $x_{i}$ is initially encoded using one-hot encoding and serves as input for both networks. The first neural network outputs $\mu_{k_{i}}$, while the second provides $\sigma^{2}_{k_{i}}$, which are then used to normalize $x_{i}$.
(a) SCB-Norm
(b) UCB-Norm
Figure 3: Visualization of Clustering Patterns in Activations: The t-SNE visualization of activations from models trained on CIFAR-10 with SCB-Norm and UCB-Norm normalization layers. The figure illustrates the formation of distinct mixture components aligning with predicted target classes, showcasing improved clustering patterns over the course of training.
Normalization Layer Description
BN Layer with batch normalizing transform (ref. Section II-A)
MN Layer with mixture normalizing transform (ref. Section II-B)
SCB-Norm Layer with supervised cluster-based normalization transform (ref. Algorithm 1)
SCB-Norm-base A softened version of SCB-Norm involves employing two distinct neural networks to independently learn $\mu$ and $\sigma$ for all mixture components. The cluster identifier one-hot encoding $onehot(k)$ serves as input for both neural networks.
UCB-Norm Layer with unsupervised cluster-based normalization (ref. Algoirthm 2)
TABLE I: Normalization method used in the experiments
Detailed descriptions of the methods under consideration are provided in Table I.
IV-A Datasets
The experiments in this study make use of several benchmark datasets that are widely recognized within the classification community:
  • CIFAR-10: A dataset with 50,000 training images and 10,000 test images, each of size $32\times 32$ pixels, distributed across 10 classes [13].
  • CIFAR-100: Derived from the Tiny Images dataset, it consists of 50,000 training images and 10,000 test images of size $32\times 32$, divided into 100 classes grouped into 20 superclasses [14].
  • Tiny ImageNet: A dataset that is a reduced version of the ImageNet dataset, containing 200 classes with 500 training images and 50 test images per class [15].
  • MNIST digits: Contains 70,000 grayscale images of size $28\times 28$ representing the 10 digits, with around 6,000 training images and 1,000 testing images per class [16].
  • SVHN: A challenging dataset with over 600,000 digit images, focusing on recognizing digits and numbers in natural scene images [17].
Dataset Task
CIFAR-10 and Tiny ImageNet This datasets are used for training a shallow convolutional neural network (ref. Section IV-B).
CIFAR-100 Three experiments (ref. Sections IV-BIV-C, and IV-D) involve the use of this dataset. These experiments encompass the training of shallow and deep convolutional neural networks, as well as the training of a generative neural network.
MNIST digits and SVHN This datasets are used to train the AdaMatch [18] model in unsupervised domain adaptation (ref. Section IV-E).
TABLE II: Summary of datasets and associated training experiments
Table II provides a visual representation of the mapping between each dataset and the specific experiments in which it is utilized.
IV-B A Comparative Study: Cluster-Based Normalization vs. Batch Normalization and Mixture Normalization Using Shallow Convolutional Neural Network
In this study, we employ a shallow convolutional neural network (Shallow CNN) structure outlined in Table III.
layer type size kernel (stride, pad)
input input 3 $\times$ 32 $\times$ 32 _ _
conv1 conv+bn+relu 64 $\times$ 32 $\times$ 32 5 $\times$ 5 (1, 2)
pool1 max pool 64 $\times$ 16 $\times$ 16 3 $\times$ 3 (2, 0)
conv2 conv+bn+relu 128 $\times$ 16 $\times$ 16 5 $\times$ 5 (1, 2)
pool2 max pool 128 $\times$ 8 $\times$ 8 3 $\times$ 3 (2, 0)
conv3 conv+bn+relu 128 $\times$ 8 $\times$ 8 5 $\times$ 5 (1, 2)
pool3 max pool 128 $\times$ 4 $\times$ 4 3 $\times$ 3 (2, 0)
conv4 conv+bn+relu 256 $\times$ 4 $\times$ 4 3 $\times$ 3 (1, 1)
pool4 avg pool 256 $\times$ 1 $\times$ 1 4 $\times$ 4 (1, 0)
linear linear 10, 100 or 200 _ _
TABLE III: Shallow CNN Architecture as Described in the Mixture Normalization (MN) Paper: A Comparison of MN and BN under Small and Large Learning Rate Regimes.
Five models are constructed based on the Shallow CNN (see Table III). The first model, denoted as BN, aligns with the foundational architecture of the Shallow CNN, incorporating only Batch Normalization layers for normalization. The remaining four models—MN, SCB-Norm, SCB-Norm-base, and UCB-Norm—are obtained by substituting the third Batch Normalization layer in the Shallow CNN with the respective layers of MN, SCB-Norm, SCB-Norm-base, and UCB-Norm (see Table I). As per the findings in [6], we conduct experiments employing two distinct learning rates, where one is five times larger than the other.
The mini-batch size is fixed at 256, and all models undergo training for 100 epochs using the AdamW optimizer [19, 20] with a weight decay of $41e-$ 4. In the training phase on CIFAR-10, CIFAR-100, and Tiny ImageNet datasets, we utilize the MN method, which entails estimating a Gaussian mixture model through Maximum Likelihood Estimation (MLE). For comparative analysis, the three mixture components identified by the Expectation-Maximization (EM) algorithm on MN serve as separate clusters (prior knowledge) ($K=3$) for SCB-Norm and SCB-Norm-base. In UCB-Norm, we define the number of mixture components for estimating as $K=3$. Our main goal is not to attain state-of-the-art results, which typically demand computationally expensive architectures and meticulous parameter tuning. Rather, our focus is on illustrating that the replacement or integration of our cluster-based normalization technique (SCB-Norm and UCB-Norm) can enhance the convergence rate, resulting in improved test accuracy. This underscores the substantial impact of our approach in enhancing model performance.
(a) BN
(b) MN
(c) SCB-Norm
(d) UCB-Norm
Figure 4: Using a randomly selected mini-batch on CIFAR-100, we depict the distribution of activations related to a subset of 128 channels within the ”conv2” layer of the Shallow CNN architecture (specified in Table III). The solid green curve represents the probability density function, while dashed curves depict distinct mixture components, each distinguished by various colors.
In Figure  4, we illustrate the activation distribution for models employing CB-Norm (SCB-Norm, UCB-Norm) and those without CB-Norm (BN, MN). Models using CB-Norm exhibit a closer resemblance to the normal distribution and display probability densities with reduced skewness. Clearly, CB-Norm provides a more accurate approximation of $p(x)$, as shown by the solid green curve, exhibiting a closer alignment with the underlying distribution compared to models without CB-Norm. Figure 7 highlights the swifter convergence of SCB-Norm, its compact variant SCB-Norm-base, and UCB-Norm in comparison to BN and MN. This accelerated convergence translates into enhanced performance on the validation dataset, with an average accuracy improvement of 2% on CIFAR-10, 3% on CIFAR-100, and 4% on Tiny ImageNet, as detailed in Table IV. This positive trend holds consistent across varying class numbers (10, 100, 200), even with an augmented learning rate from 0.001 to 0.005. The widening gap in convergence with the elevated learning rate underscores the effectiveness of our normalization technique in adeptly harnessing higher learning rates during training.
During the training of the Shallow CNN on the CIFAR-10 dataset, we consistently captured snapshots at 25-epoch intervals, integrating both SCB-Norm and UCB-Norm techniques. Following this, the trained models were employed on a randomly chosen CIFAR-10 batch, and the resulting activations underwent t-SNE visualization. The visual representations revealed the formation of clusters closely resembling the target classes predicted by the individual models. Notably, as training progressed, a noticeable refinement in these clusters was observed, contributing to a comprehensive improvement in performance, as depicted in Figure 3.
(a) CIFAR-10
(b) CIFAR-100
(c) Tiny ImageNet
Figure 5: Learning rate = 0.001
(d) CIFAR-10
(e) CIFAR-100
(f) Tiny ImageNet
Figure 6: Learning rate = 0.005
Figure 7: Test error curves when Shallow CNN (detailed in Table III) architecture is trained under different learning rate. We observe that on CIFAR-10, CIFAR-100 and Tiny ImageNet, Cluster-Based Normalization method (SCB-Norm, SCB-Norm-base and UCB-Norm) performs consistently in both small and large learning rate regimes.
CIFAR-10
model learning rate 25 epochs 50 epochs 75 epochs 100 epochs
BN-1 0.001 84,34 86,49 86,41 86,90
BN-2 0.005 83.44 84.39 85.65 85.65
MN-1 0.001 84,45 86,60 86,6 87,07
MN-2 0.005 83.47 84.60 85.68 85.80
SCB-Norm-base-1 0.001 85,9 86,93 87,09 87,11
SCB-Norm-base-2 0.005 83.65 85.09 86.28 86.41
SCB-Norm-1 0.001 84.68 86.61 87.38 87.57
SCB-Norm-2 0.005 84.85 86.54 87.07 87.07
UCB-Norm-1 0.001 85.15 85.80 87.97 87.97
UCB-Norm-2 0.005 86.04 86.04 87.49 87.49
CIFAR-100
model learning rate 25 epochs 50 epochs 75 epochs 100 epochs
BN-1 0.001 57,41 59,74 59,82 59,82
BN-2 0.005 55.64 56.97 56.29 56.29
MN-1 0.001 56,90 59,80 59,80 60,10
MN-2 0.005 55.80 55.80 56.56 57.09
SCB-Norm-base-1 0.001 57,74 61,01 61,01 61,01
SCB-Norm-base-2 0.005 55.47 57.06 57.06 57.60
SCB-Norm-1 0.001 58.50 59.25 59.65 61.02
SCB-Norm-2 0.005 55.81 57.68 57.68 57.68
UCB-Norm-1 0.001 59.44 61.10 61.10 61.10
UCB-Norm-2 0.005 58.31 60.29 60.29 60.29
Tiny ImageNet
model learning rate 25 epochs 50 epochs 75 epochs 100 epochs
BN-1 0.001 37.17 37.17 37.17 37.17
BN-2 0.005 34.11 34.11 34.11 34.11
MN-1 0.001 38,18 38,17 38,5 38,5
MN-2 0.005 34.56 34.99 35.12 35.27
SCB-Norm-base-1 0.001 38.53 40.08 40.08 40.08
SCB-Norm-base-2 0.005 34.78 35.44 35.44 36.71
SCB-Norm-1 0.001 37.72 39.45 40.29 40.29
SCB-Norm-2 0.005 37.97 38.31 38.31 38.31
UCB-Norm-1 0.001 39.71 39.71 39.71 39.71
UCB-Norm-2 0.005 35.78 37.03 37.03 37.03
TABLE IV: Evaluating Performance on CIFAR-10, CIFAR-100, and Tiny ImageNet Using the Shallow CNN architecture (see Table III) with the Integration of Batch Normalization (BN), Mixture Normalization (MN), and Cluster-Based Normalization (SCB-Norm-base, SCB-Norm and UCB-Norm). Each model is associated with a specific numerical identifier, denoting the learning rate.
To substantiate the hypothesis that CB-Norm can significantly contribute to accelerating the convergence of a neural network training, we will employ a Deep Convolutional Neural Network in the following experiment using the CIFAR-100 dataset.
IV-C A Comparative Study: Cluster-Based Normalization vs. Batch Normalization and Mixture Normalization Using Deep Convolutional Neural Network
We have demonstrated that the incorporation of cluster-based normalization techniques (SCB-Norm, SCB-Norm-base, and UCB-Norm) results in enhanced convergence speed and improved performance in models where batch normalization serves as the baseline layer normalization. It’s worth noting that our experiments were conducted on a relatively shallow architecture comprising only four convolutional layers. This raises a valid question about whether this observed behavior persists in more extensive and contemporary architectures. To investigate this, we select DenseNet [21] as our preferred architecture. We leverage two foundational architectures featuring 40 and 100 layers, aligning with the structure outlined in the mixture normalization paper. These architectures encompass three dense blocks and two transition layers, maintaining a consistent growth rate of 12. DenseNet is characterized by its dense connectivity pattern, where each layer receives direct input from all preceding layers within a dense block. This unique structure promotes feature reuse and enhances model expressiveness. All models undergo 200 epochs of training on the CIFAR-100 dataset, with a batch size of 64 and utilizing Nesterov’s accelerated gradient [22]. The learning rate initiates at $0.1$ and is reduced by a factor of 10 at $50\%$ and $75\%$ of the total training epochs. Additionally, weight decay and momentum values are set to $10^{-4}$ and $0.9$, respectively. To compare the three normalization methods (batch normalization, mixture normalization, and cluster-based normalization), we create five models for each variant (DenseNet-40 and DenseNet-100). The initial model reflects the foundational structure of DenseNet, featuring Batch Normalization layers exclusively for normalization. The second model mirrors the fundamental DenseNet architecture but introduces the MN layer as the initial layer. Models three through five correspondingly involve integrating SCB-Norm-base, SCB-Norm, and UCB-Norm layers as the initial layer in the base DenseNet model. Placing these normalization layers as the initial layer facilitates the direct infusion of cluster information for adaptive normalization directly into the image. In the case of constructing components with MN, we set $K=5$. Additionally, components derived from the EM algorithm are regarded as prior clusters ($K=5$) for SCB-Norm-base and SCB-Norm. For the unsupervised version of CB-Norm, we also adopt $K=5$ for the number of mixture components.
Figure 8: DenseNet-40
Figure 9: DenseNet-100
Figure 10: Experiments on CIFAR-100 using DenseNet [21] are showcased in Figure 10 and 10. The training/test error and cross-entropy loss are depicted for DenseNet with $40$ layers (DenseNet-40) and DenseNet with $100$ layers (DenseNet-100). Notably, the incorporation of CB-Norm (SCB-Norm-base, SCB-Norm, and UCB-Norm) significantly enhances the training process, expediting optimization. Concurrently, it demonstrates superior generalization, as evidenced by a consistently substantial margin compared to its batch normalized and mixture normalized counterparts in the error curve.
Throughout the CIFAR-100 training regimen, we systematically recorded snapshots at intervals of 25 epochs, incorporating the SCB-Norm and UCB-Norm techniques. Subsequently, the trained models were applied to a randomly selected CIFAR-100 batch, and the resulting activations underwent t-SNE visualization. The visual representations exposed the emergence of clusters closely mirroring the target classes predicted by the respective models. Notably, a discernible refinement in these clusters was observed with progressive training, contributing to a comprehensive performance enhancement, as illustrated in Figure 11.
(a) SCB-Norm
(b) UCB-Norm
Figure 11: Visualization of Clustering Patterns in Activations: The t-SNE visualization of activations from models trained on CIFAR-100 with SCB-Norm and SCB-Norm normalization layers. The figure illustrates the formation of distinct clusters aligning with predicted target classes, showcasing improved clustering patterns over the course of training.
Figure 10 vividly demonstrates that the integration of CB-Norm (SCB-Norm-base, SCB-Norm, and SCB-Norm) goes beyond simply expediting optimization during training—it consistently yields superior generalization across different architectural configurations. This advantage becomes particularly pronounced in the more intricate setting of the DenseNet architecture. Utilizing DenseNet with 40 layers (DenseNet-40) on CIFAR-100, the test error for the Batch Normalization (BN) and Mixture Normalization (MN) variants stands at 0.38 and 0.37, respectively, whereas the Cluster-Based Normalization counterpart achieves a notably lower test error of 0.36. Transitioning to the deeper DenseNet with 100 layers (DenseNet-100), the test errors for the BN and MN models drop to 0.3 and 0.29, respectively, while the cluster-based normalization variants achieve the best test error of 0.28.
In the following section, we will demonstrate a specialized application of CB-Norm with the unsupervised version (UCB-Norm) in image generation.
IV-D Unsupervised Cluster-Based Normalization (UCB-Norm) Generative Adversial Networks (GANs)
GANs, or Generative Adversarial Networks [23], are deep learning models used to generate realistic data, such as images [24, 25, 26]. The GAN architecture consists of two competing neural networks: a generator and a discriminator. The generator creates data, while the discriminator evaluates its authenticity. The training process occurs through iterations, where the generator seeks to improve its ability to deceive the discriminator, while the discriminator aims to become more effective at distinguishing generated data from real data.
However, a common challenge encountered when using GANs is the issue of ”mode collapse”. This phenomenon occurs when the generator produces only a restricted subset of possible data, leading to a loss of diversity in the generated results. In other words, the generator may focus on producing a specific type of data, neglecting the generation of other potential variations. This problem can compromise the quality and variety of the generated data, requiring specific techniques and strategies to address and enhance the overall performance of the GAN model. UCB-Norm independently standardizes internal activations across diverse disentangled modes of variation, utilizing a normalization-based approach with multiple mixture components. We postulate that integrating this approach into the generator has the potential to markedly improve the training procedure of GANs.
To test our hypothesis, we examine the Deep Convolutional Generative Adversarial Network (DCGAN) [24], which is a dedicated GAN specifically crafted for image generation tasks. In the architecture We used, the generator is composed of a linear layer followed by four deconvolutional layers. The first three deconvolutional layers are succeeded by a batch normalization layer and then a LeakyReLU [27] activation function. In this setup, the linear layer likely serves as an initial mapping from a latent space to a higher-dimensional representation. The four deconvolutional layers, also known as transposed convolutional layers, work to upsample the input progressively, transforming it into a realistic output. Batch normalization is applied after the first three deconvolutional layers, aiding in stabilizing and expediting the training process. The LeakyReLU activation function introduces non-linearity, enabling the model to learn more intricate mappings. We replace the batch normalization layers associated with the first two deconvolution layers with mixture normalization (utilizing three mixture components) and unsupervised cluster-based normalization, also with $K=3$ mixture components. The training of all models is conducted on CIFAR-100 for 200 epochs, employing the Adam optimizer [20] with $\alpha=0.0002$, $\beta_{1}=0$, and $\beta_{2}=0.9$ for both the generator and discriminator. The assessment of GAN quality is gauged using the ”Fréchet Inception Distance” (FID) [28], calculated every $10$ epochs for computational efficiency.
Figure 12: Unsupervised Cluster-Based Normalization (UCB-Norm) incorporated into the architecture of the deep Convolutional Generative Adversarial Network (DCGAN) [24]. Our findings indicate that integrating our proposed cluster-based normalization into the generator of DCGAN (DCGAN-UCB-Norm) enhances the efficiency of the training process. In contrast to the standard DCGAN employing batch normalization (DCGAN-BN) and DCGAN using mixture normalization (DCGAN-MN), DCGAN-UCB-Norm not only demonstrates accelerated convergence but also attains superior (lower) Fréchet Inception Distance (FID) scores.
Figure 12 illustrates that the DCGAN incorporating cluster-based normalization (DCGAN-UCB-Norm) exhibits not only a quicker convergence compared to its batch-normalized and mixture-normalized counterparts but also achieves superior (lower) Fréchet Inception Distance (FID) scores.
(a) DCGAN-BN
(b) DCGAN-MN
(c) DCGAN-UCB-Norm
Figure 13: Examples of generated images at epoch 200 are showcased for DCGAN-BN, DCGAN-MN, and DCGAN-UCB-Norm in Figure 13(a),13(b), and13(c), respectively.
At the conclusion of the model training, Figure 13 displays examples of generated images produced by DCGANs with batch normalization, mixture normalization, and cluster-based normalization.
In the upcoming experiments, we will focus on practical approaches to constructing prior knowledge clusters in the case of SCB-Norm without relying on any specific algorithm. This method enhances the applicability of the supervised version of CB-Norm. We will employ diverse neural network architectures to thoroughly explore the versatility. The mixture normalization, requiring data partitioning with the EM algorithm to find normalization parameters corresponding to the components, cannot be applied as a baseline.
IV-E Cluster-Based Normalization In Domain Adaptation
In this experimental setup, we showcase how the effectiveness of cluster-based normalization in improving local representations can result in substantial advancements in domain adaptation. As elucidated in [29], domain adaptation entails harnessing knowledge acquired by a model from a related domain, where there is an ample amount of labeled data, to enhance the model’s performance in a target domain characterized by limited labeled data. In this context, we consider two domains (clusters) ($K=2$): the ”source domain” and the ”target domain”. To exemplify this, we employ the two versions of cluster-based normalization (SCB-Norm with its variant SCB-Norm-base, and UCB-Norm) in conjunction with AdaMatch [18], a method that integrates the tasks of unsupervised domain adaptation (UDA), semi-supervised learning (SSL), and semi-supervised domain adaptation (SSDA). In UDA, we have access to a labeled dataset from the source domain and an unlabeled dataset from the target domain. The objective is to train a model capable of generalizing effectively to the target dataset. It’s important to note that the source and target datasets exhibit variations in distribution. Specifically, we utilize the MNIST dataset as the source dataset, while the target dataset is SVHN. Both datasets encompass various factors of variation, including texture, viewpoint, appearance, etc., and their domains, or distributions, are distinct from each other.
A model, referred to as AdaMatch [30] (using batch normalization as normalization), is trained from the ground up using wide residual networks [31] on pairs of datasets, serving as the baseline model. The training of this model involves utilizing the Adam optimizer [20] with a cosine decay schedule, gradually reducing the initial learning rate initialized at 0.03. SCB-Norm, SCB-Norm-base, and UCB-Norm are incorporated as initial layers in AdaMatch, leading to the creation of three AdaMatch models based on cluster-based normalization (AdaMatch+SCB-Norm, AdaMatch+SCB-Norm-base, and AdaMatch+UCB-Norm). This integration allows the cluster identifier (source domain and target domain) to influence the image normalization process. Since the labels in the source domain are known, the model can offer a more accurate representation of this domain in comparison to the target domain where the labels are unknown. This advantageous aspect is harnessed by considering the cluster identifier of the source domain during inference in AdaMatch+SCB-Norm and AdaMatch+SCB-Norm-base, thereby improving the model’s performance on the target domain.
MNIST (source domain)
model accuracy precision recall f1-score
AdaMatch 97.36 87.33 79.39 78.09
AdaMatch+SCB-Norm-base 99.26 99.20 99.32 99.26
AdaMatch+SCB-Norm 98.92 98.93 98.92 98.92
AdaMatch+UCB-Norm 98.9 98.5 98.90 98.95
SVHN (target domain)
model accuracy precision recall f1-score
AdaMatch 25.08 31.64 20.46 24.73
AdaMatch+SCB-Norm-base 43.10 53.83 43.10 47.46
AdaMatch+SCB-Norm 54.70 59.74 54.70 54.55
AdaMatch+UCB-Norm 33.4 43.83 40.28 42.87
TABLE V: Comparing model performance: AdaMatch vs. AdaMatch+SCB-Norm-base, AdaMatch+SCB-Norm and AdaMatch+UCB-Norm on MNIST (source) and SVHN (target) datasets.
Undoubtedly, within a wider context, Table V unequivocally highlights a remarkable enhancement in validation metrics attributable to the incorporation of cluster-based normalization. This advancement is conspicuous, manifesting as a substantial boost in accuracy by 18.02% with SCB-Norm-base, 29.62% with SCB-Norm, and 8.32%. Such marked improvement notably fortifies the performance of the AdaMatch model, leading to a significantly expedited convergence during training. The difference in performance between the supervised approach (SCB-Norm and its variant SCB-Norm-base) and the unsupervised approach (UCB-Norm) can be explained by the fact that SCB-Norm takes prior knowledge information (clusters that are source and target domains), as input and directly integrates it into the normalization process. This allows for a specific representation for each domain. In contrast, UCB-Norm does not take prior knowledge as input but only considers the number of mixture components to be estimated, which is $K=2$. Estimating mixture components in this manner can be challenging in this use case, as the information from each component is incorporated into all images, regardless of their domain. This can strongly influence the representation of the data and consequently impact performance.
(a) Source domain (MNIST)
(b) Target domain (SVHN)
Figure 14: Evolution of the gradient variance during the training of AdaMatch and AdaMatch+SCB-Norm models on the source (MNIST) and target (SVHN) domains. The figures on the left correspond to the maximum gradient variance for each epoch, while the figures on the right correspond to the average gradient variance per epoch.
Figure 14 visually underscores the invaluable contribution of cluster-based normalization through SCB-Norm in stabilizing the gradient across the entire training process, yielding benefits for both the source and target domains. Consequently, this stabilization significantly contributes to accelerated convergence and an overall enhancement in model performance.
In the following experiment, we will demonstrate another approach to construct prior knowledge (clusters) on supervised version (SCB-Norm) on a Vision Transformer [32] and then on a Convolutional Neural Network (CNN).
IV-F Supervised Cluster-Based Normalization (SCB-Norm) Using Superclasses As Prior Knowledge
As delineated in Section IV-A, both the CIFAR-100 dataset and the Oxford-IIIT Pet dataset incorporate a superclass structure alongside their class divisions. In our proposed supervised cluster-based normalization approach, we harness this superclass information as ”prior knowledge” for classification tasks. Specifically, each superclass is treated as a distinct cluster. We have incorporated the adaptive normalization technique into the Vision Transformer (ViT) architecture for CIFAR-100 classification. For our ViT implementation, we utilized a Keras baseline provided by [33]. Additionally, for the Oxford-IIIT Pet dataset, we adopted a Convolutional Neural Network (CNN) architecture as our baseline, as described in [34].
IV-F1 Using CIFAR-100 Superclass as Prior Knowledge in supervised cluster-based normalization
The central innovation in this experiment revolves around utilizing CIFAR-100 superclasses as prior knowledge (clusters) for predicting the dataset’s 100 classes, particularly in the case of SCB-Norm-base and SCB-Norm methods.
Three distinct models were employed: the base ViT model obtained from Keras [33], a modified version incorporating a Batch Normalization (BN) layer as its initial component, and two alternative models that replaced the BN layer with SCB-Norm-base and SCB-Norm layers. Training included early stopping based on validation performance, and images were pre-processed by normalizing them with respect to the dataset’s mean and standard deviation. Data augmentation techniques, such as horizontal flipping and random cropping, were applied to enhance the dataset. The AdamW optimizer, with a learning rate of $10^{-3}$ and a weight decay of $10^{-4}$, was chosen to prevent overfitting and optimize model parameters [19, 20].
model accuracy precision recall f1-score
ViT+BN 55.63 8.96 90.09 54.24
ViT+SCB-Norm-base 65.87 23.36 98.53 65.69
ViT+SCB-Norm 67.38 22.93 99.00 67.13
TABLE VI: Assessing the Performance of CIFAR-100 Using the ViT Architecture [33] incorporating BN, SCB-Norm-base, and SCB-Norm, with CIFAR-100 superclasses serving as prior knowledge or clusters.
Table VI presents the notable performance enhancements achieved by our innovative Supervised Cluster-Based Normalization (SCB-Norm and its variant SCB-Norm-base) in comparison to training the ViT architecture from scratch on CIFAR-100 with Batch Normalization (BN). The SCB-Norm approach demonstrates an accuracy improvement of approximately 12% over BN. Furthermore, it is noteworthy that both SCB-Norm-base and SCB-Norm exhibit faster convergence compared to BN, requiring less training time to attain superior performance. This observation is further supported by the comparison of train and validation loss depicted in Figure 15, indicating that SCB-Norm promotes accelerated learning while improving classification performance. These findings suggest that the proposed SCB-Norm method not only stabilizes data distributions and mitigates internal covariate shift but also significantly reduces training time for enhanced results. ViT+SCB-Norm-base and ViT+SCB-Norm achieve outstanding performance, surpassing all known ViT models when trained from scratch on the CIFAR-100 dataset.
Figure 15: Contrasting Training and Validation Loss Curves: Within the ViT Architecture on CIFAR-100, SCB-Norm and SCB-Norm-base exhibit swifter convergence and reduced validation loss, enhancing learning efficiency and classification performance compared to BN.
Figure 16: Projection of ViT+SCB-Norm outputs after applying Principal Component Analysis (PCA) for dimension reduction. Coloring is based on superclasses. A clear separation is observed between the superclasses, representing the prior knowledge (clusters). This separation leads to an enhanced representation, significantly facilitating the classification task. The green line on the histograms corresponds to the probability density function (pdf).
After training the ViT model with the SCB-Norm layer (ViT+SCB-Norm), it is applied to a randomly selected batch of data to analyze the effect of SCB-Norm on the representation of data according to prior knowledge (superclasses using as distinct clusters). After retrieving the outputs, Principal Component Analysis (PCA) is applied, with $5$ components, to visualize the behavior of our supervised cluster-based normalization method. Figure 16 illustrates this approach, showing a representation of data according to prior knowledge. The emerging cluster structure provides a better representation, thereby contributing to class separation and facilitating prediction.
IV-F2 Using Oxford-IIIT Superclass as Prior Knowledge in supervised cluster-based normalization
In this experiment, mirroring the one in the preceding section with CIFAR-100, we leverage the superclasses (dog and cat) from the Oxford-IIIT Pet Dataset to classify the 37 categories within the dataset. In this methodology, we employ a CNN (OxfordCNN) as the foundational model [34] and integrate our SCB-Norm layer just before the neural network’s prediction layer (linear layer). This experiment illustrates that supervised cluster-based normalization can function as a layer at various levels of the neural network, thereby enhancing both performance and convergence.
Model Epoch Metrics
Accuracy Precision Recall F1-Score
OxfordCNN base Epoch 80 63.12 50.23 80.11 69.23
Epoch 113 66.55 53.11 84.2 72.15
OxfordCNN+SCB-Norm-base Epoch 80 64.90 52.34 83.10 73.11
Epoch 113 67.52 55.12 86.13 75.90
OxfordCNN+SCB-Norm Epoch 80 67.92 54.24 89.11 80.12
Epoch 113 73.61 69.12 92.34 90.11
TABLE VII: Assessing Performance on the Oxford-IIIT Pet Dataset Using the OxfordCNN Architecture [34], incorporating SCB-Norm-base and SCB-Norm with ”dog” and ”cat” superclasses serving as prior knowledge (distinct clusters).
The models underwent training for 200 epochs with early stopping criteria based on validation performance. Image preprocessing involved normalization using the mean and standard deviation of the dataset. Data augmentation techniques, including horizontal flipping and random cropping, were employed to enhance the dataset. The training utilized the Adam optimizer [20] with a learning rate of $10^{-4}$.
The results presented in Table VII reinforce the earlier observations with CIFAR-100. It is apparent that cluster-based normalization leads to significantly faster convergence, a trend further supported by the performance achieved by SCB-Norm. SCB-Norm reaches a performance level after 80 epochs that surpasses the baseline (OxfordCNN), which takes 113 epochs to reach convergence.
V Analyzing Computational Complexity In Detail
The computational overhead of cluster-based normalization, when compared to batch normalization [4] and mixture normalization [6], stems from the task of estimating mixture components parameters.Let $B$ denote a batch of activations located in $\mathbb{R}^{d}$, where $d$ signifies the dimensionality of activations.
Batch normalization computes the mean ($\mu_{B}$) and standard deviation ($\sigma_{B}$) by considering the entire set of activations within the batch $B$. In terms of complexity, this involves calculating the mean ($\mu$) and standard deviation ($\sigma$) for each dimension of the batch (see Equation 1), resulting in a computational complexity of $\mathcal{O}(d)$. Additionally, this method requires learning two parameters, $\gamma$ and $\beta$, for each dimension of the batch, contributing to a parameter learning complexity of $\mathcal{O}(d)$. Therefore, the overall complexity, considering both computational and parameter aspects, is $\mathcal{O}(d)$ for the entire batch $B$.
Normalization Method Computational Complexity Parameter Complexity
Batch Normalization (BN) $\mathcal{O}(d)$ $\mathcal{O}(d)$
Mixture Normalization (MN) $\mathcal{O}(d\cdot K\cdot\log(K))$; $\mathcal{O}(d\cdot n\cdot K)$
Supervised Cluster-Based Normalization (SCB-Norm-base) - $\mathcal{O}(d)$
Supervised Cluster-Based Normalization (SCB-Norm) - $\mathcal{O}(K\cdot d)$
Unsupervised Cluster-Based Normalization (UCB-Norm) - $\mathcal{O}(K\cdot d)$
TABLE VIII: Complexities in Estimating Parameters for Normalization Methods. Computational Complexity corresponds to the time required for parameter estimation, while Parameter Complexity corresponds to the number of parameters to be estimated. Here, $d$ represents the dimension of the data, $K$ denotes the number of mixture component centers, and $n$ the number of samples.
The computational challenge associated with mixture normalization arises from the intricate process of accurately estimating the parameters of the Gaussian mixture model (GMM). This process involves a two-stage approach, where a seeding procedure is first employed to initialize the centers of the mixture components. Subsequently, the Expectation-Maximization (EM) algorithm is iteratively applied to estimate the parameters of the mixture model. In the context of the mixture normalization paper, K-means++ [35] is specifically chosen as the seeding procedure due to its applicability without assuming any characteristics about the underlying data. The complexity of the K-means++ algorithm for a given batch $B$ of samples with dimension $d$ is $\mathcal{O}(d\cdot K\cdot\log(k))$, where $K$ is the number of cluster center. For the EM algorithm, the complexity of each iteration is $\mathcal{O}(d\cdot n\cdot K)$, where $n$ is the number of samples in $B$. The EM algorithm typically converges in a small number of iterations, often less than $10$ or $20$. Consequently, the overall complexity of running K-means++ followed by the EM algorithm on a batch of samples $B$ is primarily determined by the initialization step, resulting in $\mathcal{O}(d\cdot K\cdot\log(K))$, with an additional contribution from the EM iterations. The total number of iterations for the EM algorithm depends on the convergence criteria and the specific characteristics of the data. In practice, the overall complexity is often dominated by the initialization step for K-means++.
For our cluster-based normalization approach, we will detail the complexity for the supervised version (SCB-Norm and its variant SCB-Norm-base) and the unsupervised version (UCB-Norm). Consider the activation batch $B$, with dimension $d$. Assuming that the defined number of prior knowledge (clusters) is $K$, the complexity of SCB-Norm is correlated with $K$. Specifically, for each cluster $k$, two learnable parameters of dimension $d$ are associated with it ($\mu_{k}$ and $\sigma_{k}$). Therefore, the total number of parameters to be learned for all clusters is $2\times K\times d$, resulting in a complexity of $\mathcal{O}(K\cdot d)$ for SCB-Norm. SCB-Norm-base, being a lighter variant, utilizes two linear layers of dimension $d$. It employs these two linear layers across all clusters $K$ to estimate the parameters $\mu$ and $\sigma$, reducing the complexity from $\mathcal{O}(K\cdot d)$ to $\mathcal{O}(d)$. The unsupervised version, UCB-Norm, estimates three learnable parameters for each mixture component $K$: $\lambda_{k}$ of dimension 1, and $\mu_{k}$ and $\sigma_{k}$ of dimension $d$. This results in a total of $2\times K\times d+K$ parameters to be learned, with a complexity of $\mathcal{O}(K\cdot d)$. Unlike mixture normalization, both the supervised and unsupervised versions of cluster-based normalization are performed in a one-stage process, and the complexity is solely associated with the parameter estimation during the backpropagation process. A detailed comparison is presented in Table VIII.
VI Conclusion
In conclusion, Cluster-Based Normalization (CB-Norm) emerges as a groundbreaking one-step normalization method, addressing challenges in deep neural network training. Its adaptability and unique approach, incorporating Gaussian mixture models, mark a significant advancement in overcoming issues like internal covariate shift and gradient stability. Through Supervised Cluster-Based Normalization (SCB-Norm) and Unsupervised Cluster-Based Normalization (UCB-Norm), CB-Norm demonstrates versatility in handling diverse learning scenarios. SCB-Norm utilizes predefined data partitioning (clusters), while UCB-Norm dynamically clusters activations witout using prior knowledge, both contributing to enhanced adaptability. Our experiments aim to assess CB-Norm’s performance, comparing it with Batch Normalization (BN) and Mixture Normalization (MN). We apply SCB-Norm and UCB-Norm to various tasks, including classification, domain adaptation, and generative neural networks. Furthermore, we introduce SCB-base, a simplified variant, maintaining a fixed cluster identifier ($k$) and leveraging neural networks to encode cluster information. These experiments collectively aim to showcase the efficacy and versatility of CB-Norm in improving deep neural network performance across a spectrum of applications.